# Noise-Tolerant Hybrid Approach for Data-Driven Predictive Control

### v1

@@ 4.2 Noisy Case: Robustness to Measurement Noise
$V(k)=\mathbb{E}[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$
In order to study positive definiteness of $V(k)$, under Assumption 4.1, we have $l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$ so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|)\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|),$ for some $\alpha_{1,W}\in\mathcal{K}_{\infty}$. For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)+\sigma_{y}^{*}(k)$ and bounded sets $\mathcal{Y},\mathcal{U}$ (Assumption 4.8) imply that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$ and noise effects. By Assumption 4.8:
$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}.$
Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
Combining, there exist $\alpha_{1,V},\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that:
$\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
As dissipation inequality, we construct a suboptimal solution at $k+1$ like (53) with $\sigma_{y,s}(k+1)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k+1)+\mathbf{P}_{2}\mathbf{u}_{s}(k+1)-\mathbf{y}_{s}(k+1)$, where $\hat{u}(N|k),\hat{y}(N|k)$ satisfy Assumption 4.7. The suboptimal cost is:
$J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}.$
By computing $\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]$ as follows:
$\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (58)
Since:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (59)
Then, by subtracting (58) and (59), we have:
$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]$ $=-\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]-\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
$\quad+\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]+\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
By Assumption 4.7:
$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(id-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (60)
For the slack terms, by Assumption 4.6:
$\|\sigma_{y,s}(k+1)\|\leq\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k+1)-\mathbf{z}_{\text{ini}}(k+1))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k+1)\|$ (61)
so, by taking the expectation from both side of (61):
$\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k+1)\|^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\bar{c}.$
Similarly, $\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}$. Let $c_{\sigma,\text{max}}=\max(c_{\sigma},\lambda_{\max}(\Lambda_{y})L_{P}^{2})$. Then:
$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$ (62)
And for $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (63)
By taking the expectation from both side of (63), thus:
$\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]=\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (64)
By combining (62) and (64):
$V(k+1)-V(k)\leq\mathbb{E}[J_{s}(k+1)-J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$ (65)
By substituting (60) in (65), yields:
$V(k+1)-V(k)\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
By Assumption 4.1, there exists $\alpha_{4}\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]\geq\alpha_{4}(\mathbb{E}[\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|])$ (66)
Since $\tilde{X}_{\text{ini}}(k)$ includes these terms and using (66), there exists $\alpha_{4,V}\in\mathcal{K}_{\infty}$ such that:
$V(k+1)-V(k)\leq-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
The term $\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]$ is bounded by the predictor’s output and noise effects, contributing to the ISS gain.
To show ISS property, By Assumption 4.5, $\bar{c}\leq c_{\zeta}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2}$, the dissipation inequality:
$V(k+1)\leq V(k)-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\alpha_{5}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+c_{\zeta,\text{max}}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2},$
where $\alpha_{5}\in\mathcal{K}_{\infty}$ bounds the control cost, and $c_{\zeta,\text{max}}=2c_{\sigma,\text{max}}c_{\zeta}$. Applying stochastic ISS results [25], there exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
Since $V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}$ and $\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]\geq 0$:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c},$
ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state plus a noise-dependent term. Hence, the proof is completed. ∎
Remark 4.3.
The ISS property ensures bounded tracking errors under noise, as validated in Fig. 2, where NTDPC maintains stable tracking of $(r_{y},r_{u})$ with $\sigma_{\zeta}^{2}=0.25I$. The sensitivity index $I_{s}$ (Fig. 5) supports Assumption 4.8, aiding horizon selection.
@@ 5 Numerical Simulations
$\begin{split}&U:=\left\{u\in\mathbb{R}^{2}:\begin{bmatrix}-20\\ -20\end{bmatrix}\leq u\leq\begin{bmatrix}20\\ 20\end{bmatrix}\right\},\quad\\ &Y:=\left\{y\in\mathbb{R}^{2}:\begin{bmatrix}-25\\ -15\end{bmatrix}\leq y\leq\begin{bmatrix}25\\ 15\end{bmatrix}\right\}.\end{split}$ (68)
@@ 5.1 Noise-free case
The Fig.1 illustrates the performance of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$) when tracking constant admissible references ($(r_{y},r_{u})$), as analyzed in Section 5.1. The results are consistent with the asymptotic stability proven in Theorem 4.4, where $y(k)\to r_{y}$ and $u(k)\to r_{u}$ as $k\to\infty$ under Assumptions 4.1, 4.2, and 4.3. The top plot displays the primary output $y_{1}(k)$ for NTDPC (blue), SPC (red), and SMMPC (green) over time, with the dashed line representing the reference $r_{y}$; all methods converge to $r_{y}$, but NTDPC exhibits the fastest transient response and minimal overshoot, reflecting the effectiveness of the data-driven predictor and terminal condition. The second plot shows a secondary output $y_{2}(k)$, with similar convergence behavior, where NTDPC (blue) stabilizes more quickly than SPC (red) and SMMPC (green), supporting the robustness of the NTDPC framework to initial conditions encoded in $\mathbf{z}_{\text{ini}}(k)$. The third plot presents the primary input $u_{1}(k)$, where NTDPC (blue) adjusts rapidly from an initial value, stabilizing near $r_{u}$ (dashed line) with less oscillation compared to SPC (red) and SMMPC (green), consistent with the bounded input constraints in $\mathcal{U}$. Finally, the bottom plot depicts a secondary input $u_{2}(k)$, showing NTDPC (blue) achieving a steady state closer to $r_{u}$ with reduced variability, further validating the stability properties derived from the Lyapunov function $V(\mathbf{z}_{\text{ini}}(k))$. Overall, the convergence of outputs and inputs to their respective references demonstrates the asymptotic stability of NTDPC, as predicted by the storage function analysis $V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))$, and the superior performance of NTDPC over SPC and SMMPC highlights the advantage of the noise-tolerant predictor and optimization framework, particularly in the absence of noise.
@@ 5.2 Noisy case
To quantify controller performance, the objective function is evaluated and plotted in Fig. 3 which reveals the superiority of NTDPC and SPC compared to SMMPC. It is worthy to note that there is a slight difference between NTDPC and SPC throughout the entire simulation, highlighting the effectiveness of the designed NTDPC. In summary, the figures provide a comprehensive comparison of NTDPC, SPC and SMMPC control strategies, highlighting their performance in terms of output and input trajectories and their respective variability. This comparison aids in understanding the strengths and weaknesses of each method in maintaining desired control objectives amidst system noise and constraints.

### v2

@@ 4.2 Noisy Case: Robustness to Measurement Noise
$\mathbb{E}[\Delta_{\text{stage}}\mid\mathcal{F}_{k}]\leq L_{l}\mathbb{E}[\|\mathbf{y}(k)-\mathbf{y}^{*}(0|k)\|\mid\mathcal{F}_{k}]\leq L_{l}\mathbb{E}[\|\mathbf{\zeta}(k)\|\mid\mathcal{F}_{k}]\leq L_{l}\sqrt{\bar{c}}.$ (67)
Combining these results:
$\mathbb{E}[V(k+1)-V(k)\mid\mathcal{F}_{k}]\leq-\rho(\alpha_{4,V}(\mathbb{E}[\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|]))+\bar{C}_{\text{total}},$ (68)
where $\bar{C}_{\text{total}}$ aggregates all noise-dependent constants. This inequality confirms that the Lyapunov function decreases outside a bounded region determined by the noise level, which is the definition of ISS [23].
@@ 5 Numerical Simulations
[…]
@@ 5.1 Noise-free case
The Fig.2 illustrates the performance of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$) when tracking constant admissible references ($(r_{y},r_{u})$), as analyzed in Section 5.1. The results are consistent with the asymptotic stability proven in Theorem 4.4, where $y(k)\to r_{y}$ and $u(k)\to r_{u}$ as $k\to\infty$ under Assumptions 4.1, 4.2, and 4.3. The top plot displays the primary output $y_{1}(k)$ for NTDPC (black), SPC (red), and SMMPC (green) over time, with the dashed line representing the reference $r_{y}$; all methods converge to $r_{y}$, but NTDPC exhibits the fastest transient response and minimal overshoot, reflecting the effectiveness of the data-driven predictor and terminal condition. The second plot shows a secondary output $y_{2}(k)$, with similar convergence behavior, where NTDPC (black) stabilizes more quickly than SPC (red) and SMMPC (green), supporting the robustness of the NTDPC framework to initial conditions encoded in $\mathbf{z}_{\text{ini}}(k)$. The third plot presents the primary input $u_{1}(k)$, where NTDPC (black) adjusts rapidly from an initial value, stabilizing near $r_{u}$ (dashed line) with less oscillation compared to SPC (red) and SMMPC (green), consistent with the bounded input constraints in $\mathcal{U}$. Finally, the bottom plot depicts a secondary input $u_{2}(k)$, showing NTDPC (black) achieving a steady state closer to $r_{u}$ with reduced variability, further validating the stability properties derived from the Lyapunov function $V(\mathbf{z}_{\text{ini}}(k))$. Overall, the convergence of outputs and inputs to their respective references demonstrates the asymptotic stability of NTDPC, as predicted by the storage function analysis $V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))$, and the superior performance of NTDPC over SPC and SMMPC highlights the advantage of the noise-tolerant predictor and optimization framework, particularly in the absence of noise.
@@ 5.2 Noisy case
To quantify controller performance, the objective function is evaluated and plotted in Fig. 4 which reveals the superiority of NTDPC and SPC compared to SMMPC. It is worth noting that there is a slight difference between NTDPC and SPC throughout the entire simulation, highlighting the effectiveness of the designed NTDPC. In summary, the figures provide a comprehensive comparison of NTDPC, SPC and SMMPC control strategies, highlighting their performance in terms of output and input trajectories and their respective variability. This comparison aids in understanding the strengths and weaknesses of each method in maintaining desired control objectives amidst system noise and constraints.